Wilmington Company has two manufacturing departments—Assembly and Fabrication. It considers all of its manufacturing overhead costs to be fixed costs. The first set of data that is shown below is based on estimates from the beginning of the year. The second set of data relates to one particular job completed during the year—Job Bravo.

Estimated Data Assembly Fabrication Total
Manufacturing overhead costs $ 1,400,000 $ 1,680,000 $ 3,080,000
Direct labor-hours 70,000 42,000 112,000
Machine-hours 28,000 140,000 168,000
Job Bravo Assembly Fabrication Total
Direct labor-hours 15 7 22
Machine-hours 7 10 17

Required:

1. If Wilmington used a plantwide predetermined overhead rate based on direct labor-hours, how much manufacturing overhead would be applied to Job Bravo?

2. If Wilmington uses departmental predetermined overhead rates with direct labor-hours as the allocation base in Assembly and machine-hours as the allocation base in Fabrication, how much manufacturing overhead would be applied to Job Bravo?

(Round your intermediate calculation to 2 decimal places.)

1. Plantwide manufacturing overhead applied to Job Bravo
2. Manufacturing overhead applied from Assembly to Job Bravo
Manufacturing overhead applied from Fabrication to Job Bravo
Total departmental manufacturing overhead applied to Job Bravo $0
Answer #1

1. Estimated MOH = $3,080,000

Estimated DLH = 112,000

Predetermined OH rate = Estimated MOH/Estimated DLH

= $3,080,000/112,000 = $27.5/DLH

Plant wide MOH applied to Job Bravo = Actual DLH * Predetermined rate per DLH

=22 * 27.5 = $605

2. Assembly department OH rate = Estimated MOH in Assembly dept / Estimated DLH in Assembly dept

=$1,400,000/70,000 = $20/ DLH

MOH applied from Assembly to Job Bravo = Actual DLH in Assembly dept * Assembly dept OH rate

= 15*20 =$300

Fabrication department OH rate =Estimated MOH in fabrication dept / Estimated machine hours in Fabrication dept

= $1,680,000/140,000

=$12/MH

MOH applied from fabrication to Job bravo = Actual MH in fabrication dept * Fabrication dept OH rate

= 10*12 = $120

1 Plant wide MOH applied to Job Bravo $605
2 MOH applied from Assembly to Job Bravo $300
MOH applied from fabrication to Job Bravo $120
Total departmental MOH applied to Job Bravo $1025